OpenStudy (anonymous):

(FOG)(X) f(x)= x^2+2x and g(x)= 1/x

OpenStudy (lgbasallote):

do you mean (f o g)(x)?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

that means you substitute g(x) into EVERY x of f(x)...so what is your g(x)?

OpenStudy (anonymous):

1/x

OpenStudy (anonymous):

(1/x)^2 + 2(1/x)

OpenStudy (lgbasallote):

that's right..simplify

OpenStudy (anonymous):

1/(x^2) + 2/x ?

OpenStudy (lgbasallote):

good there you go...are you tasked to combine it?

OpenStudy (anonymous):

nope.. Don't think so it just says "find (f o g)(x)?"

OpenStudy (lgbasallote):

well that's your answer then...if you want you can combine it to be safe..use x^2 as LCD \[\frac{1}{x} + \frac{2x}{x^2} = \frac{1 + 2x}{x^2}\]

OpenStudy (lgbasallote):

the first term is supposed to be \(\large \frac{1}{x^2}\)

OpenStudy (anonymous):

Thank you!